OpenStudy (anonymous):

The graph of g(x) is f(x) translated to the left 8 units and up 2 units. What is the function rule for g(x) given f(x) = x²?

OpenStudy (anonymous):

af(k(x-d))+c kd affects the x ac affects the y

OpenStudy (klimenkov):

If a function \(y=f(x)\) is translated to the left \(a\) units and up \(b\) units then it must transform in this way: \(y=f(x-a)+b\)

OpenStudy (anonymous):

Okay thanks that makes since - just one more question... Is the "²" inside the parenthesis (sp?) or outside of them?

OpenStudy (klimenkov):

OOps, i made a mistake. \(y=f(x+a)+b\)

OpenStudy (anonymous):

Ughh the negative was good...?